# Dynamical critical behavior on the Nishimori point of frustrated Ising models

### v1

@@ I Introduction
Critical phenomena have been a paramount concept of statistical physics, which involves many interesting features such as universality Stanley (1971); Goldenfeld (2018); Cardy (1996). The text book example of such a system is the Ising model, where the low temperature ferromagnetic phase is separated from a high temperature paramagnetic phase via a second order phase transition point. The critical exponents in the pure Ising system are universal with respect to microscopic details — a property which is strongly endorsed by the renormalization group (RG) theory and many numerical works Stanley (1971); Goldenfeld (2018); Cardy (1996); Privman and Fisher (1984); Hu et al. (1995). The universality concept is not limited to static properties as one can also identify dynamical universality classes Hohenberg and Halperin (1977); Täuber (2014); Ódor (2004); Wang and Hu (1997) which are influenced by the conservation laws imposed dynamically. In second order phase transitions the latter are classified according to the values taken by the dynamic critical exponents. This idea came from the application of RG methods to dynamical properties Halperin et al. (1972); Hohenberg and Halperin (1977); Halperin and Hohenberg (1969); Suzuki (1977). Two systems having the same set of static critical exponents can possess different dynamical exponents if, for example, they evolve with non-conserved or conserved order parameter dynamics. Therefore, in order to distinguish the universality class of a system, one needs both the static and dynamical critical exponents.
@@
Here, $S_{i}=\pm 1$ are Ising spins, placed at each site $i$ of a $d=2$ lattice. The subscript $\langle ij\rangle$ denotes the sum over all nearest neighbor (nn) spin pairs, and the exchange interaction couplings $J_{ij}$ are quenched random variables drawn from a probability distribution $P(J_{ij})$ (see below). A negative value of $J_{ij}$ indicates an anti-ferromagnetic bond (which introduces frustration), while a positive value represents a ferromagnetic bond. To examine the universality, we focus on two types of bond distributions. One is the bimodal (±$\pm J$) law
@@
[…]
@@ II Numerical techniques and definitions
SL $T_{\rm N}\simeq 0.953$ $T_{\rm N}\simeq 0.959$
TL $T_{\rm N}\simeq 1.228$ $T_{\rm N}\simeq 1.253$
HL $T_{\rm N}\simeq 0.759$ $T_{\rm N}\simeq 0.787$
Table 1: Values of the critical temperature $T_{\rm N}$ at the Nishimori point (NP) in the $2d$ frustrated Ising model on the square lattice (SL), triangular lattice (TL), and honeycomb lattice (HL), with bimodal (±$\pm J$) and Gaussian $(G)$ bond distributions. These values are quoted from earlier studies de Queiroz (2009, 2006); Ohzeki (2009); Merz and Chalker (2002) and references therein. The frustration parameter ($p_{\rm N}$, or $r_{\rm N}$) can be calculated using the relations in Eq. (25).
@@
When a spin system is thermally quenched to a critical point, the initial state becomes unstable and the system prefers to reach the new critical equilibrium. In order to minimize the net free energy, the spins locally group as in the target critical state — resulting in the growth of (spatial) correlations with time. To investigate these correlations quantitatively, the standard practice is to consider the two-point space-time correlation function
@@
Figure 1: Plot of the correlation length $\xi(t)$ vs time $t$, in log-log scale, for critical Ising quenches, that is to $T_{\rm Is}$, of systems with different lattice geometries (see the key) from infinitely high temperature $T\gg T_{\rm Is}$ initial conditions. The linear size of each system is fixed to $L=512$. The dashed line indicates the pure growth law $\xi(t)\sim t^{1/z_{\rm c}}$, with $z_{\rm c}\sim 2.17$.
@@ III Main results
system $z_{\rm c}$ $\lambda_{\rm c}/z_{\rm c}$
SL, ±$\pm J$ $6.31(5)$ $0.224(3)$
SL, $G$ $6.22(3)$ $0.21(1)$
TL, ±$\pm J$ $6.39(4)$ $0.219(4)$
TL, $G$ $6.49(3)$ $0.217(3)$
HL, ±$\pm J$ $6.18(5)$ $0.222(5)$
HL, $G$ $6.35(4)$ $0.224(2)$
Table 2: Values of the exponents $z_{\rm c}$ and $\lambda_{\rm c}/z_{\rm c}$, for a quench to the Nishimori point (NP) at $T=T_{\rm N}$ from a high temperature ($T\gg T_{\rm N}$) state, in the $2d$ frustrated Ising model defined on different geometries [square lattice (SL), triangular lattice (TL), and honeycomb lattice (HL)] with different bond distributions [bimodal (±$\pm J$) and Gaussian ($G$)]. The values and corresponding (statistical) error bars are estimated from power-law fits to the long time behavior of the quantities $\xi(t)$ and $q_{0}(t)$.
@@ III.1 Growth of the dynamical correlation length
Now, let us discuss the results for a quench to the NP ($T=T_{\rm N}$) from an infinitely high temperature $T\gg T_{\rm N}$. In the main frame of Fig. 2, the correlation length $\xi(t)$ is plotted against $t$ for systems with different geometries and bond distributions. The linear size of each system is again $L=512$, with a non-equilibrium ensemble average performed over $1000$ independent initial configurations and disorder realizations. In the asymptotic regime, all the datasets exhibit power law behavior of the form $\xi(t)\sim t^{1/z_{\rm c}}$, on the log-log scale of the plot. To estimate the value of the exponent $z_{\rm c}$ from a dataset, we performed various power-law fits by varying the time-window within the temporal regime $t\in[10^{4},10^{6}]$. The fits with the values of reduced-$\chi^{2}$ near $1$ are accepted, and the corresponding estimates of the exponent $z_{\rm c}$ are mentioned in Table 2. The statistical error bars are obtained using Jackknife resampling technique Efron (1982), which also accounts for the correlations in a dataset.
We notice that apart from the statistical errors reported in Table 2 the fits may also have systematic errors, which are often harder to capture. For this purpose, we first employ different ways of calculating $\xi(t)$. For all the systems reported in Table 2 the exponent $z_{\rm c}$ varies significantly beyond the statistical error bars; however, we always obtained $z_{\rm c}$ between $6.0-6.5$. Furthermore, the localization of the NP also ensues the errors in $z_{\rm c}$, as the values of $T_{\rm N}$ reported in Table 1 are just numerical estimates. To account for these errors, we vary the value of $T_{\rm N}$ by about $0.01$. We find again that the exponent $z_{\rm c}$ changes but remains within the range $6.0-6.5$.
@@
In the inset of Fig. 2, $z_{\rm eff}(t)$ is plotted against $t$ for different datasets in the main frame. The asymptotic value of $z_{\rm eff}(t)$ for different lattices and bond distributions fluctuates between $6.0-6.5$. However, the data in the inset also show that the fluctuations are of oscillatory nature, which eventually should die out in a longer simulation.
@@ III.2 Critical aging
Now let us understand the aging phenomenon during the critical dynamics at the NP. We first present the results for a quench from the paramagnetic state ($T\gg T_{\rm N}$) to $T=T_{\rm N}$. This set up follows the zero-field-cooling (ZFC) protocol, but with the initial temperature $T=\infty$.
@@
Figure 5: Main frame: Log-log plot of the autocorrelation function $C(t,t_{w})$ against time difference $t-t_{w}$ of a system defined on the square lattice (SL) with bimodal (±$(\pm J)$ disorder, for a quench from the Ising critical point $T_{\rm Is}$ to the Nishimori point $T_{\rm N}$. Different datasets correspond to waiting times $t_{w}\in[10,10000]$. Inset: Plot of $q_{0}(t)$ vs $t$, also in log-log scale, for a quench from $T_{\rm Is}$ to $T_{\rm N}$ of SL systems with ±$\pm J$ and Gaussian ($G$) bond distributions (see the key). The dashed lines in both plots denote the expected asymptotic behavior $f(t)\sim t^{-\eta/z_{c}}$, with $\eta=0.18$ and $z_{c}=6$.
@@ III.3 Global persistence
Table 3: Values of exponent $\theta_{\rm c}z_{\rm c}$, for a quench from a sharply prepared initial state (with magnetization density $M(0)=0.004$) to the Nishimori Point (NP) at $T=T_{\rm N}$, in the $2d$ frustrated Ising model defined on different geometries [square lattice (SL), triangular lattice (TL), and honeycomb lattice (HL)] with bimodal (±$\pm J$) distribution. Different values are obtained from the scaling collapse in Fig. 7. See the main text for details.
@@
Let us now discuss the universality features of the global persistence probability $p(t)$ (13) after a quench from high $T$ to the NP. In Fig. 6, we first present data for $p(t)$ vs $t$ of an initial condition with $\langle M(0)\rangle=0$. The idea is that the magnetization for randomly generated configurations of a small system will likely attain a nonzero value, with ensemble averaged magnetization $\langle M(0)\rangle=0$. The linear size of the system is $L=128$, with an average over around $10^{5}$ nonequilibrium runs. For large $t$, all the datasets belonging to different lattices and bond disorders in the present figure exhibit power-law decay, $p(t)\sim t^{-\theta_{\rm c}}$, with $\theta_{\rm c}\simeq 0.080~{}(05)$.
@@
Figure 7: Plot of the scaling variables $L^{\theta_{c}z_{\rm c}}p(t)$ vs $t/L^{z_{\rm c}}$, in log-log scale, for a quench to the Nishimori point $T_{\rm N}$ from a sharply prepared initial states with nonzero magnetization density ($M(0)=0.004$). Different datasets represent systems of different lattice geometries with ±$\pm J$ quenched-disorder (see the key) and different linear sizes: the linear size of the system for the datasets with empty symbols is $L=128$, while for the datasets with filled symbols it is $L=64$.
@@
To gain some precision for the evaluation of $\theta_{\rm c}$, we consider the quench dynamics of a sharply prepared initial state, with magnetization density $M(0)=0.004$. To get rid of the finite size effects, the expression in Eq. (14) can be rewritten using a scaling relation da Silva et al. (2003); Majumdar et al. (1996) as
$p(t)=L^{-\theta_{\rm c}z_{\rm c}}f\left(\frac{t}{L^{z_{\rm c}}}\right)\;.$ (19)
@@
To check the above relation, the scaling variable $L^{\theta_{c}z_{\rm c}}p(t)$ is plotted against the scaled time $t/L^{z_{\rm c}}$ in Fig. 7. Datasets represent systems of linear sizes $L=64$ and $L=128$, each defined on different lattice geometries with ±$\pm J$ disorder (the Gaussian case is not shown for brevity). The nonequilibrium averages are performed over $5\times 10^{5}$ and $10^{5}$ runs for $L=64$ and $L=128$, respectively. The values of $\theta_{\rm c}z_{\rm c}$ in the scaling variable $L^{\theta_{c}z_{\rm c}}p(t)$ are chosen so as to obtain data-collapse, whereas $z_{\rm c}$ is fixed to $z_{\rm c}\simeq 6$ in the time-axis. After preasymptotic time scales, a good scaling collapse is obtained for all datasets. Even for early time scales, a collapse is observed for each system size. The obtained values of exponent $\theta_{\rm c}z_{\rm c}$ are listed in Table 3. Upon the increase in $L$, the value of $\theta_{\rm c}z_{\rm c}$ seems to decrease for systems with different geometries. When using $z_{c}=6$, a value of $\theta_{\rm c}$ between $0.07-0.08$ is obtained, which is also in agreement with $\theta_{\rm c}\simeq 0.080\pm 0.005$ observed in Fig. 6. Finally, we conclude that the persistence exponent attains a universal value around $\theta_{\rm c}=0.080\pm 0.005$.

### v2

@@ I Introduction
Critical phenomena have been a paramount concept of statistical physics, which involves many interesting features such as universality Stanley (1971); Goldenfeld (2018); Cardy (1996). The text book example of such a system is the Ising model, where the low temperature ferromagnetic phase is separated from a high temperature paramagnetic phase via a second order phase transition point. The critical exponents in the pure Ising system are universal with respect to microscopic details — a property which is strongly endorsed by the renormalization group (RG) theory and many numerical works Stanley (1971); Goldenfeld (2018); Cardy (1996); Privman and Fisher (1984); Hu et al. (1995). The universality concept is not limited to static properties as one can also identify dynamical universality classes Hohenberg and Halperin (1977); Täuber (2014); Ódor (2004); Wang and Hu (1997) which are influenced by the conservation laws imposed dynamically. In second order phase transitions the latter are classified according to the values taken by the dynamic critical exponents. The idea of dynamical universality came from the application of RG methods to dynamical properties Halperin et al. (1972); Hohenberg and Halperin (1977); Halperin and Hohenberg (1969); Suzuki (1977). Two systems having the same set of static critical exponents can possess different dynamical exponents if, for example, they evolve with non-conserved or conserved order parameter dynamics. Therefore, in order to distinguish the universality class of a system, one needs both the static and dynamical critical exponents.
@@
Here, $S_{i}=\pm 1$ are Ising spins, placed at each site $i$ of a $2d$ lattice. The subscript $\langle ij\rangle$ denotes the sum over all nearest neighbor (nn) spin pairs, and the exchange interaction couplings $J_{ij}$ are quenched random variables drawn from a probability distribution $P(J_{ij})$ (see below). A negative value of $J_{ij}$ indicates an anti-ferromagnetic bond (which introduces frustration), while a positive value represents a ferromagnetic bond. To examine the universality, we focus on two types of bond distributions. One is the bimodal (±$\pm J$) law
@@
Due to various intriguing features like the existence of nontrivial fixed points of different universality classes including a multicritical point and emergence of a zero-temperature spin glass phase, the model (1) is quite popular in the literature Dotsenko and Dotsenko (1982); McMillan (1984); Le Doussal and Harris (1988); Honecker et al. (2001); Picco et al. (2006); Merz and Chalker (2002); Amoruso and Hartmann (2004); Hasenbusch et al. (2008); Parisen Toldin et al. (2009); Ohzeki (2009); Agrawal et al. (2023a). It is also exploited to determine the error correction threshold for certain quantum error correction codes Kitaev (1997, 2003); Iba (1999); Dennis et al. (2002). The key property which makes this model so important is the existence of a peculiar line in the phase diagram — the Nishimori line (see Refs. Nishimori (2001, 1980, 1981); Le Doussal and Harris (1988, 1989); Gruzberg et al. (2001); Krzakala and Zdeborová (2011); Garban and Spencer (2022) and Appendix A for further details). Le Doussal and Harris Le Doussal and Harris (1988, 1989) found that the multicritical point of the paramagnetic - ferromagnetic (PF) boundary is located on the intersection of the Nishimori line with the PF boundary; therefore, it is also known as the Nishimori point (NP). In previous studies Honecker et al. (2001); de Queiroz (2009), the central charge $c$ at NP was numerically calculated, $c\simeq 0.46$. We also recently Agrawal et al. (2023a) found that the interfaces at this multicritical point can be uniquely described by a Stochastic Loewner Evolution (SLEκ) with $\kappa\simeq 2.2$. The description of the critical interfaces in terms of SLEκ consolidates the conformal invariance of the system; also see Refs. Amoruso et al. (2006); Bernard et al. (2007); Picco and Santachiara (2008) for the SLEκ description of $2d$ spin glasses and nonminimal Conformal Field Theories (CFTs).
@@ II Numerical techniques and definitions
      SL       $T_{\rm N}=0.953(2)$       $T_{\rm N}=0.958(1)$
      TL       $T_{\rm N}=1.228(1)$       $T_{\rm N}=1.254(2)$
      HL       $T_{\rm N}=0.759(1)$       $T_{\rm N}=0.788(1)$
Table 1: Values of the critical temperature $T_{\rm N}$ at the Nishimori point (NP) in the $2d$ frustrated Ising model on the square lattice (SL), triangular lattice (TL), and honeycomb lattice (HL), with bimodal (±$\pm J$) and Gaussian $(G)$ bond distributions. These values are quoted from earlier studies de Queiroz (2009, 2006); Ohzeki (2009); Merz and Chalker (2002) and references therein. The corresponding error bars account for the range of $T_{\rm N}$ considered for the localization estimates. The frustration parameter ($p_{\rm N}$, or $r_{\rm N}$) can be calculated using the relations in Eq. (25).
@@
When a spin system is thermally quenched to a critical point, the initial state becomes unstable and the system prefers to reach the new critical equilibrium. In order to minimize the net free energy, the spins locally group as in the target critical state — resulting in the growth of (spatial) correlations with time. To investigate these correlations quantitatively, the standard practice is to consider the two-point spatial correlation function
@@
Figure 1: Plot of the correlation length $\xi(t)$ vs time $t$, in log-log scale, for critical Ising quenches, that is to $T_{\rm Is}$, of systems with different lattice geometries (see the key) from infinitely high temperature $T\gg T_{\rm Is}$ initial conditions. The linear size of each system is fixed to $L=512$. The dashed line indicates the pure growth law $\xi(t)\sim t^{1/z_{\rm c}}$, with $z_{\rm c}=2.17$.
@@ III Main results
      system       $z_{\rm c}$       $\lambda_{\rm c}/z_{\rm c}$       $\theta_{\rm c}$
      SL, ±$\pm J$       $6.31(5)$       $0.224(3)$       $0.083(2)$
      SL, $G$       $6.22(3)$       $0.21(1)$       $0.077(5)$
      TL, ±$\pm J$       $6.39(4)$       $0.219(4)$       $0.082(3)$
      TL, $G$       $6.49(3)$       $0.217(3)$       $0.081(2)$
      HL, ±$\pm J$       $6.18(5)$       $0.222(5)$       $0.084(2)$
      HL, $G$       $6.35(4)$       $0.224(2)$       $0.085(1)$
Table 2: Values of the exponents $z_{\rm c}$, $\lambda_{\rm c}/z_{\rm c}$ and $\theta_{\rm c}$, for a quench to the Nishimori point (NP) at $T=T_{\rm N}$ from a high temperature ($T\gg T_{\rm N}$) state, in the $2d$ frustrated Ising model defined on different geometries [square lattice (SL), triangular lattice (TL), and honeycomb lattice (HL)] with different bond distributions [bimodal (±$\pm J$) and Gaussian ($G$)]. The values and corresponding (statistical) error bars are estimated from the best power-law fits to the long time behavior of the quantities $\xi(t)$, $q_{0}(t)$ and $p(t)$, with the reduced $\chi^{2}$-value being near $1$.
@@ III.1 Growth of the dynamical correlation length
Now, let us discuss the results for a quench to the NP ($T=T_{\rm N}$) from an infinitely high temperature $T\gg T_{\rm N}$. In the main frame of Fig. 2, the correlation length $\xi(t)$ is plotted against $t$ for systems with different geometries and bond distributions. The linear size of each system is again $L=512$, with a non-equilibrium ensemble average performed over $1000$ independent initial configurations and disorder realizations.
In the late time regime, all the datasets exhibit power law behavior of the form $\xi(t)\sim t^{1/z_{\rm c}}$ over a three decades of time, on the log-log scale of the plot. Such a promising behavior spread over multiple decades should correspond to the asymptotic limit. To estimate the value of the exponent $z_{\rm c}$ from a dataset, we performed various power-law fits by varying the time-window within the temporal regime $t\in[10^{4},10^{6}]$. The fits with the values of reduced-$\chi^{2}$ near $1$ are accepted, and the corresponding estimates of the exponent $z_{\rm c}$ are mentioned in Table 2. The statistical error bars are obtained using Jackknife resampling technique Efron (1982), which also accounts for the correlations in a dataset. However, we do not take these error bars seriously as they do not account for the systematic errors explained below.
Apart from the statistical errors reported in Table 2 the fits may also have systematic errors, which are often harder to capture. For this purpose, we first employ different ways of calculating $\xi(t)$. For all the systems reported in Table 2 the exponent $z_{\rm c}$ varies significantly beyond the statistical error bars; however, we always obtained $z_{\rm c}$ between $6.0-6.5$. Furthermore, the localization of the NP also ensues the errors in $z_{\rm c}$, as the values of $T_{\rm N}$ reported in Table 1 are just numerical estimates. To account for these errors, we vary the value of $T_{\rm N}$ by about $0.01$. We find again that the exponent $z_{\rm c}$ changes but remains within the range $6.0-6.5$.
@@
In the inset of Fig. 2, $z_{\rm eff}(t)$ is plotted against $t$ for different datasets in the main frame. The asymptotic value of $z_{\rm eff}(t)$ for different lattices and bond distributions fluctuates between $6.0-6.5$. However, the data in the inset also show that the fluctuations are of oscillatory nature, which eventually should die out in a longer simulation. Furthermore, for an independent check of the correlation length $\xi(t)$, we analyze the fluctuations $\Delta M(t)=\overline{\langle M^{2}\rangle}-\overline{\langle M\rangle}^{2}$ in the magnetization density $M(t)$. According to the short-time critical dynamics (STCD) approach Tomé and de Oliveira (1998); *Albano_2011, this quantity behaves as $\Delta M(t)\sim t^{\zeta}$, with $\zeta=(d-\eta)/z_{c}$ after a quench from infinitely high temperature. This analysis gives a value of $z_{\rm c}$ between $6.1-6.6$ for the different systems considered in this work (data not shown here).
@@ III.2 Critical aging
Now let us understand the aging phenomenon during the critical dynamics on the NP. We first present the results for a quench from the paramagnetic state ($T\gg T_{\rm N}$) to $T=T_{\rm N}$. This set up follows the zero-field-cooling (ZFC) protocol, but with the initial temperature $T=\infty$.
@@
Figure 5: Main frame: Log-log plot of the autocorrelation function $C(t,t_{w})$ against time difference $t-t_{w}$ of a system defined on the square lattice (SL) with bimodal (±$(\pm J)$ disorder, for a quench from the Ising critical point $T_{\rm Is}$ to the Nishimori point $T_{\rm N}$. Different datasets correspond to waiting times $t_{w}\in[10,10000]$, where $t_{w}$ increases from bottom to top. Inset: Plot of $q_{0}(t)$ vs $t$, also in log-log scale, for a quench from $T_{\rm Is}$ to $T_{\rm N}$ of SL systems with ±$\pm J$ and Gaussian ($G$) bond distributions (see the key). The dashed lines in both plots denote the expected asymptotic behavior $f(t)\sim t^{-\eta/z_{c}}$, with $\eta=0.18$ and $z_{c}=6$.
@@ III.3 Global persistence
Table 3: Values of exponent $a_{\rm c}=\theta_{\rm c}z_{\rm c}$, for a quench from a sharply prepared initial state (with magnetization density $M(0)=0.004$) to the Nishimori Point (NP) at $T=T_{\rm N}$, in the $2d$ frustrated Ising model defined on different geometries [square lattice (SL), triangular lattice (TL), and honeycomb lattice (HL)] with bimodal (±$\pm J$) distribution. Different values are obtained by enabling the scaling collapse in Fig. 7 for the two systems of linear sizes $L=64$ and $128$ defined on each geometry (individual column). The corresponding error bars denote the range of the scaling exponent $a_{\rm c}$ (for fixed $z_{\rm c}=6$) in which the collapse of the data remains statistically good.
@@
Let us now discuss the universality features of the global persistence probability $p(t)$ (13) after a quench from high $T$ to the NP. In Fig. 6, we first present data for $p(t)$ vs $t$ of an initial condition with $\langle M(0)\rangle=0$. The idea is that the magnetization for randomly generated configurations of a small system will likely attain a nonzero value, with ensemble averaged magnetization $\langle M(0)\rangle=0$. The linear size of the system is $L=128$, with an average over around $10^{5}$ nonequilibrium runs. For large $t$, all the datasets belonging to different lattices and bond disorders in the present figure are consistent with power-law decay, $p(t)\sim t^{-\theta_{\rm c}}$, with $\theta_{\rm c}\simeq 0.080~{}(05)$. The obtained values from the best power-law fits are listed in Table 2.
@@
Figure 7: Plot of the scaling variables $L^{a_{\rm c}}p(t)$ vs $t/L^{z_{\rm c}}$, in log-log scale, for a quench to the Nishimori point $T_{\rm N}$ from sharply prepared initial states with nonzero magnetization density ($M(0)=0.004$). Different datasets represent systems of different lattice geometries with ±$\pm J$ quenched-disorder (see the key) and different linear sizes: the linear size of the system for the datasets with empty symbols is $L=128$, while for the datasets with filled symbols it is $L=64$.
@@
To check what happens to the universality of $\theta_{\rm c}$ when an initial configuration with magnetization density $M(0)\neq 0$ is considered, we also investigate the quench dynamics of a sharply prepared initial state. Since $M(0)$ is tiny (we consider $M(0)=0.004$), finite size effects may eliminate its impact. Therefore, we use a scaling relation da Silva et al. (2003); Majumdar et al. (1996) obtained by rewriting Eq. (14) as,
$p(t)=L^{-a_{\rm c}}f\left(\frac{t}{L^{z_{\rm c}}}\right)\;,$ (19)
@@
where exponent $a_{\rm c}=\theta_{\rm c}z_{\rm c}$. In Fig. 7, the scaling variable $L^{a_{\rm c}}p(t)$ is plotted against the scaled time $t/L^{z_{\rm c}}$. Datasets represent systems of linear sizes $L=64$ and $L=128$, each defined on different lattice geometries with ±$\pm J$ disorder (the Gaussian case is not shown for brevity). The nonequilibrium averages are performed over $5\times 10^{5}$ and $10^{5}$ runs for $L=64$ and $L=128$, respectively. The values of $a_{\rm c}$ in the scaling variable $L^{a_{\rm c}}p(t)$ are chosen so as to obtain data-collapse, whereas $z_{\rm c}$ is fixed to $z_{\rm c}=6$ on the time-axis. After preasymptotic time scales, a good scaling collapse is obtained for datasets of both $L$. In fact, the same scaling relation (19) is obeyed by all systems of different geometries. The obtained values of exponent $a_{\rm c}$ are listed in Table 3. Upon the increase in $L$, the value of $a_{\rm c}$ seems to decrease. When using $z_{c}=6$, a value of $\theta_{\rm c}$ between $0.07-0.08$ is obtained from $a_{\rm c}=\theta_{\rm c}z_{\rm c}$, which is also in agreement with $\theta_{\rm c}\simeq 0.080\pm 0.005$ observed in Fig. 6. Finally, we conclude that the persistence exponent attains a universal value around $\theta_{\rm c}=0.080\pm 0.005$.